# Sampling reduced density matrix to extract fine levels of entanglement spectrum and restore entanglement Hamiltonian

### v5

@@ Results
Figure 1: (a) A path integral configuration of a reduced density matrix. The length of imaginary time is $\beta$, and the time boundary of $B$/$A$ is periodic/open; (b) Lattice of the Spin-1/2 Heisenberg ladder from an overhead view, with spins in $A$ and $B$ depicted by red and blue circles respectively; (c) Its phase diagram by denoting $J_{\text{leg}}=\cos\theta$ and $J_{\text{rung}}=\sin\theta$.
@@
Figure 2: (a) The comparison of low lying entanglement excitation spectra obtained by pure ED and QMC+ED with different Monte Carlo steps. The red lines are the results obtained by pure ED and the black dots are the results obtained by QMC+ED. Here we choose $L=8$. (b) Entanglement excitation spectra versus total momenta $k$ in the chain direction for $\theta=\pi/3$.
Figure 3: (a) Extrapolation of velocity $v$ as a function of $1/L$. The inset shows the fitting of velocity $v_{L}$ based on the numerical results for various size $L$. (b) Spectral function of the system with size $L=24$ and its corresponding velocity $v$ fitting. (c) Entanglement spectrum versus total momenta $k$ in the chain direction for $\theta=2\pi/3$ with the system size $L=20$. The red line is the quadratic fitting curve in the entanglement spectrum, given by $\omega=0.73k^{2}$. (d) The entanglement spectral function of FM Heisenberg ladder with $L=20$. The white line is the quadratic fitting curve near the gapless point.
@@
Figure 4: (a) The AFM Heisenberg model on the square lattice. The dashed lines are used to illustrate the bipartition into two subsystems. The yellow dashed lines illustrate the cutting method that $A$ is a ring and the red dashed lines illustrate the method that $A^{\prime}$ is a block. For each cutting method, the other part is denoted as $B$. We consider the square lattice model with size $20\times 20$. We show the entanglement spectrum of EH corresponding to (b) $A$ with length $L=20$ and (c) $A^{\prime}$ with size $4\times 4$. The TOS levels are connected by a red line. All the data are calculated in the total $S^{z}=0$ sector.
@@
Meanwhile, if we set $A$ as a $4\times 4$ region as the red dots ($A^{\prime}$ region) in Fig. 4(a), the ES also shows a TOS character as shown in Fig.4(c). Usually, the TOS holds in cornerless cutting cases, this result reveals that the TOS also arises in cornered cutting. For this $L=4$ square region, we obtain that the velocity of the TOS fitting line is about $0.38$. According to the Eq.(3), we gain the $\chi_{\perp}\sim 0.08$. The $\chi_{\perp}$ here is basically consistent with the QMC result ($\sim 0.07$ [79]) of square lattice AFM Heisenberg model in thermaldynamic limit obtained through size extrapolation  [^1] Although the ES should not be exactly same as the energy spectrum, they are always similar according to the Li-Haldane conjecture..
@@
We simulate the AFM Heisenberg ladder at $J_{\mathrm{rung}}=50,~{}2,~{}0.5$ with fixed $J_{\mathrm{leg}}=1$ to obtain the RDM and then restore the related couplings according to the above scheme. The estimates for $J_{1}$ and $J_{2}$ are displayed in the Table 1. The case with a large $J_{\mathrm{rung}}=50$ fits well with the perturbation, while $J_{\mathrm{rung}}=2$ also provides a good match. However, when $J_{\mathrm{rung}}=0.5<J_{\mathrm{leg}}=1$, the perturbation theory no longer hold. In such a case, we find that $|J_{2}|$ goes much closer to $|J_{1}|$, indicating the emergence of a long-range interaction in the EH, as shown in Refs. [85, 61]. And the coupling decays much more slowly with distance.
@@ Discussion
To compare with ED, this method can obtain $\rho_{A}$ for much larger system sizes. To compare with the QMC+SAC scheme, it is not restricted to the cut-geometry between $A$ and $B$, and it can extract the fine levels of low-lying ES  [^2] In the QMC+SAC method, because the spectral function is gained from the imaginary time correlation function via numerical analytic continuation, it requires that $A$ should have periodic boundary condition to guarantee a well-defined momentum $k$. In other words, the analytic continuation depends closely on the cut form of the entanglement. However, our method can study any geometry for the $A$.. Similarly, although the Bisognano-Wichmann theorem gives the detailed form of entanglement Hamiltonian, it holds under some strict conditions (e.g., the entanglement region has no corner, the original Hamiltonian satisfies translation symmetry) and may lose the effectiveness in many lattice models [89, 90, 91, 92].
@@
The data that support the findings of this study are available from the authors upon reasonable request.
@@
All numerical codes in this paper are available upon reasonable request to the authors.

### v6

@@ Results
Figure 1: The schematic diagram of sampling reduced density matrix and the characters of Heisenberg ladder. (a) A path integral configuration of a reduced density matrix. The length of imaginary time is $\beta$, and the time boundary of $B$/$A$ is periodic/open; (b) Lattice of the Spin-1/2 Heisenberg ladder from an overhead view, with spins in $A$ and $B$ depicted by red and blue circles respectively; (c) Its phase diagram by denoting $J_{\text{leg}}=\cos\theta$ and $J_{\text{rung}}=\sin\theta$.
@@
Figure 2: Comparison of entanglement spectra using different methods. (a) The comparison of low lying entanglement excitation spectra obtained by pure ED and QMC+ED with different Monte Carlo steps. The red lines are the results obtained by pure ED and the black dots are the results obtained by QMC+ED. Here we choose $L=8$. (b) Entanglement excitation spectra versus total momenta $k$ in the chain direction for $\theta=\pi/3$.
Figure 3: Extrapolation and spectral analysis of entanglement spectrum in the spin ladder system. (a) Extrapolation of velocity $v$ as a function of $1/L$. The inset shows the fitting of velocity $v_{L}$ based on the numerical results for various size $L$. (b) Spectral function of the system with size $L=24$ and its corresponding velocity $v$ fitting. (c) Entanglement spectrum versus total momenta $k$ in the chain direction for $\theta=2\pi/3$ with the system size $L=20$. The red line is the quadratic fitting curve in the entanglement spectrum, given by $\omega=0.73k^{2}$. (d) The entanglement spectral function of FM Heisenberg ladder with $L=20$. The white line is the quadratic fitting curve near the gapless point.
@@
Figure 4: The entanglement spectrum of entanglement Hamiltonian in 2D system. (a) The AFM Heisenberg model on the square lattice. The dashed lines are used to illustrate the bipartition into two subsystems. The yellow dashed lines illustrate the cutting method that $A$ is a ring and the red dashed lines illustrate the method that $A^{\prime}$ is a block. For each cutting method, the other part is denoted as $B$. We consider the square lattice model with size $20\times 20$. We show the entanglement spectrum of EH corresponding to (b) $A$ with length $L=20$ and (c) $A^{\prime}$ with size $4\times 4$. The TOS levels are connected by a red line. All the data are calculated in the total $S^{z}=0$ sector.
@@
Meanwhile, if we set $A$ as a $4\times 4$ region as the red dots ($A^{\prime}$ region) in Fig. 4(a), the ES also shows a TOS character as shown in Fig.4(c). Usually, the TOS holds in cornerless cutting cases, this result reveals that the cornered cutting affects TOS little, at least, in the Heisenberg model. For this $L=4$ square region, we obtain that the velocity of the TOS fitting line is about $0.38$. According to the Eq.(3), we gain the $\chi_{\perp}\sim 0.08$. The $\chi_{\perp}$ here is basically consistent with the QMC result ($\sim 0.07$ [79]) of square lattice AFM Heisenberg model in thermaldynamic limit obtained through size extrapolation. (Although the ES should not be exactly same as the energy spectrum, they are always similar according to the Li-Haldane-Poilblanc conjecture.)
@@
We simulate the AFM Heisenberg ladder at $J_{\mathrm{rung}}=50,~{}2,~{}0.5$ with fixed $J_{\mathrm{leg}}=1$ to obtain the RDM and then restore the related couplings according to the above scheme. The estimates for $J_{1}$ and $J_{2}$ are displayed in the Table 1. The case with a large $J_{\mathrm{rung}}=50$ fits well with the perturbation, while $J_{\mathrm{rung}}=2$ also provides a good match. However, when $J_{\mathrm{rung}}=0.5<J_{\mathrm{leg}}=1$, the perturbation theory no longer holds. In such a case, we find that $|J_{2}|$ goes much closer to $|J_{1}|$, indicating the emergence of a long-range interaction in the EH, as shown in Refs. [84, 61]. At the same time, the coupling decays much more slowly with distance.
@@ Discussion
To compare with ED, this method can obtain $\rho_{A}$ for much larger system sizes. To compare with the QMC+SAC scheme, it is not restricted to the cut-geometry between $A$ and $B$, and it can extract the fine levels of low-lying ES. (In the QMC+SAC method, because the spectral function is gained from the imaginary time correlation function via numerical analytic continuation, it requires that $A$ should have periodic boundary condition to guarantee a well-defined momentum $k$.) In other words, the analytic continuation depends closely on the cut form of the entanglement. However, our method can study any geometry for the $A$. Similarly, although the Bisognano-Wichmann theorem gives the detailed form of entanglement Hamiltonian, it holds under some strict conditions (e.g., the entanglement region has no corner, the original Hamiltonian satisfies translation symmetry) and may lose the effectiveness in many lattice models [87, 88, 89, 90].
@@
Methods
We use quantum Monte Carlo algorithm combined with exact diagonalization technique in this work. Details have been explained in the main text.
@@ Methods
The data that support the findings of this study are available at https://doi.org/10.5281/zenodo.14879371
@@
All numerical codes in this paper are available from the authors.